A certain reaction is second order in N2  and second order in H2. Use this information to complete the table below. Round each of your answers to 3 significant digits.

[N2] [H2] initial rate of reaction
0.311 M 2.02 M 5.00 x 104M/s
0.311 M 9.15 M _________M/s
0.884 M 0.710 M __________M/s
Answer #1

Given,

[N2] = 0.311 M

[H2] = 2.02 M

initial rate = 5 * 104 M/s

The reaction is of second order with each of the reactant, so

rate = k[N2]2[H2]2

(5 * 104 M/s) = k (0.311 M)2 (2.02 M)2

k = (5 * 104) / 0.3946

k = 12.67 * 104 M-3s-1

Given,

[N2] = 0.311 M

[H2] = 9.15 M

k = 12.67 * 104 M-3s-1

rate = k[N2]2[H2]2

rate = 12.67 * 104 * (0.311)2 (9.15)2

= 1.026 * 106 M/s

Given,

[N2] = 0.884 M

[H2] = 0.710 M

k = 12.67 * 104 M-3s-1

Now,

rate = k[N2]2[H2]2

rate = 12.67 * 104 * (0.884)2 * (0.710)2

= 4.99 * 104 M/s

The table is,

[N2] [H2] Initial rate of reaction
0.311 M 2.02 M

5.00 * 104 M/s

0.311 M 9.15 M

1.026 * 106 M/s

0.884 M 0.710 M

4.99 * 104 M/s